OpenStudy (anonymous):

Solve these two equations : 2sin^2 x - 5sin x - 3 = 0 (cos x - (square root of 2)/2)(sec x - 1) = 0

OpenStudy (anonymous):

for the first one, let u = sin(x) you now have to solve the quadratic: 2u^2 - 5u - 3 = 0 when you solve for u, replace it back with sinx

OpenStudy (anonymous):

2*-3=-6 6-1=5 6*-1=-6 \[2 \sin ^2x-\left( 6-1 \right)\sin x-3=0,~or~2~\sin ^2x-6\sin x+\sin x-3=0\] make factors and solve

OpenStudy (jdoe0001):

as Euler271 and surjithayer suggested \(\bf 2{\color{blue}{ sin}}^2{\color{blue}{ (x)}} - 5{\color{blue}{ sin(x)}} - 3 = 0\qquad {\color{blue}{ u=sin(x)}} \\ \quad \\ 2{\color{blue}{ u}}^2-5{\color{blue}{ u}}-3=0\)

OpenStudy (anonymous):

so factor like 2x^2 - 5x - 3?

OpenStudy (anonymous):

\[2 \sin x \left( \sin x-3 \right)+1\left( \sin x-3 \right)=0\] \[\left( \sin x-3 \right)\left( 2 \sin x+1 \right)=0\] either sin x-3=0, sin x=3,Rejected,because \[\left| \sin x \right|\le 1\] \[\sin x=-\frac{ 1 }{2 }=-\sin 60=\sin \left( 180+60 \right),\sin \left( 360-60 \right)\] calculate x

OpenStudy (jdoe0001):

yes, as surjithayer said already, keep in mind that \(\bf -1\le sin(x) \le 1\)

OpenStudy (anonymous):

ok now I've got that I've got to factor 2x^2 - 5x - 3

OpenStudy (anonymous):

(x - 3) (2x+1)

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

x would represent sin^2x? or just sin x? in the factorization

OpenStudy (anonymous):

anyone there?

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

x=sin x